Find the equation of the plane which passes through the line of intersection of the planes 3π‘₯ + 4𝑦 βˆ’ 5𝑧 = 9 and 2π‘₯ + 6𝑦 + 6𝑧 = 7 and which is perpendicular to the plane 3π‘₯ + 2𝑦 βˆ’ 5𝑧 + 6 = 0


1
Expert's answer
2022-02-08T15:10:46-0500

Any plane though the line of intersection of the planes:

"3𝑥 + 4𝑦 − 5𝑧 - 9=0 \text{ and } 2𝑥 + 6𝑦 + 6𝑧 - 7=0"

is given by;

"(3𝑥 + 4𝑦 − 5𝑧-9)+\lambda(2𝑥 + 6𝑦 + 6𝑧-7)=0\\
\Rightarrow x(3+2\lambda)+y(4+6\lambda)+z(-5+6\lambda)=(3+2\lambda,\ 4+6\lambda,\ -5+6\lambda)\cdot(x,\ y,\ z)=9+7\lambda\cdots\cdots(1)"and this is perpendicular to the plane "\displaystyle
3𝑥 + 2𝑦 − 5𝑧 =(3,\ 2,\ -5)\cdot(x,\ y,\ z)= -6" given in the question.


Since they are perpendicular, then their dot product of the vectors is zero. That is;

"\displaystyle
(3+2\lambda,\ 4+6\lambda,\ -5+6\lambda)\cdot(3,\ 2,\ -5)=3\times(3+2\lambda)+2\times(4+6\lambda)-5\times(-5+6\lambda)=0\\
\Rightarrow9+6\lambda+8+12\lambda+25-30\lambda=0\\
\Rightarrow\lambda=\frac{7}{2}"

Substituting "\displaystyle
\lambda=\frac{7}{2}" into "(1)" yields;

"\displaystyle
x\left[3+2\left(\frac{7}{2}\right)\right]+y\left[4+6\left(\frac{7}{2}\right)\right]+z\left[-5+6\left(\frac{7}{2}\right)\right]=9+7\left(\frac{7}{2}\right)"

"\displaystyle
\Rightarrow20x+50y+32z=67" which is the equation of the plane we want.